OpenStudy (anonymous):

Write the equation of the line which passes through (–4, 2) and is parallel to y = –x + 6 in slope-intercept form.

OpenStudy (anonymous):

Parallel lines have the same slope. So the slope of the new line is -1. Now use point-slope form to write the equation. \[y - 2 = -1(x + 4)\] Solve for y \[y = -x - 2\]